The equation of the parabola C is y = - 2x^2 + 5x.The equation of the line l is y - 4x + 1 = 0.
Fina the intercept of C with the axes.

Answer by stanbon(75887)

The equation of the parabola C is y = - 2x^2 + 5x.
Find the intercept of C with the axes.
------------------
y-intercept:: ?
Let x = 0, then y = -2*0^2+5*0 = 0
So, y-intercept is (0,0)
-----
x-intercept:: ?
Let y = 0, then -2x^2+5x = 0
Factor to get:
(x(-2x+5) = 0
x = 0 or x = 5/2
x-intercepts are (0,0) and (5/2,0)
====================================
Cheers,
Stan H.